OpenStudy (anonymous):

Simplify the following expression: 2x - 2y + 5z - 2x - y + 3z x - 3y + 8z 2x - 2y + 8z 3y + 8z -3y + 8z

OpenStudy (anonymous):

i tried working it out but i got none of the answers above

OpenStudy (texaschic101):

2x - 2y + 5z - 2x - y + 3z combine like terms 2x - 2x = 0 -2y - y = -3y 5z + 3z = 8z put it all together and you get : -3y + 8z

OpenStudy (anonymous):

thank you for the steps shown i pretty much fail at math

OpenStudy (texaschic101):

glad I could help :)